# Large-Scale Data Selection for Instruction Tuning

### v1

@@ Abstract
Selecting high-quality training data from a larger pool is a crucial step when instruction-tuning language models, as carefully curated datasets often produce models that outperform those trained on much larger, noisier datasets. Automated data selection approaches for instruction-tuning are typically tested by selecting small datasets (roughly 10k samples) from small pools (100-200k samples). However, popular deployed instruction-tuned models often train on hundreds of thousands to millions of samples, subsampled from even larger data pools. We present a systematic study of how well data selection methods scale to these settings, selecting up to 2.5M samples from pools of up to 5.8M samples and evaluating across 7 diverse tasks. We show that many recently proposed methods fall short of random selection in this setting (while using more compute), and even decline in performance when given access to larger pools of data to select over. However, we find that a variant of representation-based data selection (RDS+), which uses weighted mean pooling of pretrained LM hidden states, consistently outperforms more complex methods across all settings tested – all whilst being more compute-efficient. Our findings highlight that the scaling properties of proposed automated selection methods should be more closely examined. We release our code, data, and models at https://github.com/hamishivi/automated-instruction-selection.
@@
Hamish Ivison1,2, Muru Zhang1,3, Faeze Brahman2, Pang Wei Koh1,2, Pradeep Dasigi2 1University of Washington, 2Allen Institute for AI, 3University of Southern California hamishiv@cs.washington.edu
@@ 1 Introduction
We test nine different data selection approaches in our setting, covering a variety of different approaches to data selection: gradient-based influence methods (Xia et al., 2024), embedding-based methods (Zhang et al., 2018; Hanawa et al., 2021; Ivison et al., 2023a), and loss/perplexity-based methods (Yin and Rush, 2024; Antonello et al., 2021; Marion et al., 2023; Ankner et al., 2024; Li et al., 2024b). Surprisingly (and contrary to prior work), we find that a simple embedding method — using the hidden states of a pretrained LM — works best overall across all settings tested, consistently beating all other approaches. In particular, we find that:
@@
    Similarly, when selecting for multiple tasks at once, using RDS+ outperforms baselines at varied data sizes, selecting datasets from 10k to 2.5M samples in size. Additionally, we find that RDS+ even outperforms Tülu 2, which was trained on a human-curated mixture (§4.2).
@@
    Controlling for the compute used during selection, our RDS variant outperforms random selection by an average of two points when selecting hundreds of thousands to millions of samples, whilst using less compute. This stands in contrast to prior work (Xia et al., 2024), which does not scale up to large enough sizes to observe the point at which RDS becomes more compute-efficient. This highlights that data selection techniques may only become particularly useful for larger-scale settings than those typically used in prior work. (§4.3)
Overall, by examining performance when selecting over much larger data pools that are tied to real instances of popular instruction-tuned models, we believe that our findings better reflect how data selection approaches can aid LM training when working at the data scales used to train deployed models.
@@ 2 Related Work
Earlier approaches to data selection for instruction tuning relied primarily on careful human curation (Zhou et al., 2023) or extensive ablation studies to determine optimal data mixtures (Wang et al., 2023). Recently, many automated data selection approaches have emerged, ranging from using GPT-based quality assessment  (Chen et al., 2024; Wettig et al., 2024; Lu et al., 2023; Liu et al., 2024), embedding or perplexity features from model inference (Marion et al., 2023; Ivison et al., 2023a; Li et al., 2024b; Zhang et al., 2024c; Sachdeva et al., 2024; Li et al., 2024a; Du et al., 2023; Zhang et al., 2024b), gradient-based methods (Xia et al., 2024; Han et al., 2020; Zhang et al., 2024a; Yu et al., 2024), or shallower heuristics (Zhao et al., 2024a; Cao et al., 2024; Li et al., 2024d). These works typically focus on finding small subsets of larger datasets that can match or outperform using the entire dataset, often selecting subsets of roughly 10k samples in size (Chen et al., 2024; Xia et al., 2024; Li et al., 2024b) from datasets with 100s of thousands of samples. This does not match the known sizes of various instruction-tuning datasets used in practice, which range from 300k to millions of samples in size (Touvron et al., 2023; Llama Team, 2024; Ivison et al., 2023b; Lambert et al., 2024). In this work, we aim specifically to study how these methods perform when selecting large datasets from large pools, and find that many proposed methods fall short of random selection and an RDS baseline in this setting.
@@ 3.2 Data Pool
Figure 2: Size and makeup of data pools considered in this work (unfiltered Tulu 2, 3) and in past work (Xia et al., 2024; Chen et al., 2024; Li et al., 2024b). We provide the size of each pool on top of each bar. Each color represents a different dataset. See App. B for more details on data pool composition.
@@ 4.1 Single-Task Data Selection
Method MMLU GSM8k BBH TydiQA Codex Squad AlpacaEval Average Selecting 10k datapoints for single tasks from 200k data points Random (unbal.) 45.8 17.2 41.5 52.7 27.0 83.5 18.4 40.9 Random (bal.) 45.4 16.6 40.7 50.7 27.9 80.6 33.2 42.2 LESS 46.7 21.5 41.8 57.5 19.6 84.7 37.9 44.2 Embed (NV) 46.9 19.9 42.4 52.3 25.0 88.3 24.0 42.7 Embed (GTR) 47.1 20.2 42.4 52.4 27.7 88.3 16.7 42.1 Top-PPL 46.4 9.9 33.8 46.9 20.3 77.6 4.3 34.2 Mid-PPL 46.9 15.5 39.7 53.6 25.7 81.6 12.9 39.4 IFD 42.6 17.5 39.1 38.7 29.1 57.9 45.9 38.7 RDS+ 47.0 20.9 42.7 52.3 31.8 88.2 42.3 46.4 Selecting 10k datapoints for single tasks from all 5.8M data points Random (unbal.) 46.9 17.4 42.1 53.2 26.8 83.6 15.1 40.7 Random (bal.) 45.3 16.5 40.2 50.9 28.2 78.0 29.0 41.1 Top-PPL 44.6 6.1 23.9 40.4 20.9 74.4 2.5 30.4 Mid-PPL 46.1 14.9 40.7 52.7 23.0 82.3 12.5 38.9 Embed (GTR) 45.0 29.9 42.8 45.5 29.7 88.3 43.8 46.4 Embed (NV) 47.0 23.1 41.1 51.2 29.1 88.8 10.9 41.6 IFD 41.9 13.0 37.9 40.0 26.4 46.6 44.6 35.8 RDS+ 47.5 33.9 42.9 54.9 32.4 88.5 53.5 50.5
@@ 4.2 Multi-task Selection
Method MMLU GSM8k BBH TydiQA Codex Squad AlpacaEval Average Rand. (unbal) 52.2 18.0 44.5 55.3 25.7 81.5 33.9 44.5 Rand. (bal) 51.5 21.8 45.1 50.7 32.2 87.9 43.2 47.5 Top-PPL 49.1 10.5 39.4 49.4 21.6 80.3 5.6 36.6 Mid-PPL 53.1 13.3 42.8 52.4 20.3 86.2 20.7 41.3 Embed (GTR) 49.9 32.8 44.6 54.4 30.4 88.4 35.7 48.0 Embed (NV) 50.6 28.7 44.4 56.0 30.4 89.1 17.9 45.3 IFD 45.7 21.8 41.2 39.5 27.7 17.0 57.4 35.7 Tülu 2 50.0 22.7 45.1 54.0 33.1 86.9 54.4 49.5 RDS+ 50.2 35.2 44.7 56.3 35.1 89.0 45.6 50.9 RDS+ - Wildchat 50.9 24.8 43.6 57.3 31.1 87.3 39.3 47.8 RDS+ - Arena Hard 48.1 36.2 43.9 51.8 31.8 81.3 59.4 50.4
@@
We explore using WildChat (Zhao et al., 2024b) and Arena Hard (Li et al., 2024c) as out of distribution query sets for selecting data. Crucially, this means we do not assume access to any query samples from our test tasks. We find that using Arena Hard samples performs close to RDS+, showing we can select high-quality samples without assuming any data from the evaluations in our test suite. This suggests that using a high-quality selection set results in a data mixture that generalizes well to unseen tasks.
@@ 4.3 Scaling Multi-task Selection
(Fig. 4). This shows that targeted selection is even useful when selecting datasets 100s of thousands to millions of data points in size. Furthermore, RDS+ performs similarly to training on all data (50.8 vs 50.9) even when selecting only 326,000 samples (roughly 6% of the overall pool), and outperforms training on all data when selecting over 1M samples.
@@
(Fig. 3). When taking into account the compute used for selection and training, we find that the extra compute incurred by RDS does not pay off until we select relatively large datasets (roughly $\geq 326$k), after which it becomes significantly better than random selection. This highlights once again that examining data selection performance at scale is important to show the potential of some methods. Additionally, we note that the cost of RDS+ could potentially be further reduced by re-using existing embeddings or using smaller models for selection.
@@ 5 Analysis
We examine qualitatively what samples get selected by RDS+, IFD, and PPL (see App. H for a visualization). For RDS+, We find that the sources selected vary a reasonable amount depending on downstream evaluation, often with intuitive explanations: e.g., GSM8k requires strong chain-of-thought reasoning abilities, and so the chain-of-thought data is upweighted relative to the normal distribution. Similarly, for HumanEval-Codex the Code Alpaca dataset is upweighted.
@@ 6 Conclusion
In this work, we have explored how well a variety of data selection techniques work when selecting datasets of up to 2.5M samples from pools comprising of close to 6M data points. By examining data selection techniques in these settings, we find that many methods proposed in the past not only underperform random selection baselines, but also perform worse with larger pools of data. Embedding-based methods, and in particular RDS+, are the primary exception, outperforming all other data selection methods across various selection sizes. Finally, we find that using smaller models to select data can also work well, albeit not as well as larger ones. Exploring how to use smaller proxy models for RDS+ selection at scale is a promising future direction to further reduce the cost of automated data selection. We believe that our results overall highlight the importance of testing data selection methods over large, diverse pools of data, and testing how well data selection approaches scale with respect to both data and compute. We will publicly release our data and code to aid in future work.
Acknowledgments
We thank Hannaneh Hajishirzi, as well as members of UW NLP and Ai2, for useful advice and feedback on earlier versions of this project. We thank Mengzhou Xia for useful advice around running the LESS codebase. This project is supported by the Singapore National Research Foundation and the National AI Group in the Singapore Ministry of Digital Development and Information under the AI Visiting Professorship Programme (award number AIVP-2024-001), and by the AI2050 program at Schmidt Sciences.
@@ Appendix B Data Pool Sources Breakdown
Source # Samples in Tülu 2 # Samples in Unfil.
FLAN V2 (Chung et al., 2022) 49,123 961,322
FLAN CoT (Chung et al., 2022) 49,747 398,439
Open Assist. (Köpf et al., 2023) 7,331 7,707
Dolly (Databricks, 2023) 0 15,007
GPT-4 Alpaca (Peng et al., 2023) 19,906 52,002
Code Alpaca (Chaudhary, 2023) 20,016 20,022
ShareGPT 111,912 100,054
LIMA (Zhou et al., 2023) 1,018 1,030
Wizard Evol-Instruct V2 (Xu et al., 2023) 29,810 142,802
Open Orca (Lian et al., 2023) 29,683 4,111,858
SciRIFF (Wadden et al., 2024) 7,468 7,535
Hardcoded 140 14
Total 326,153 5,817,792
Table 5: Number of samples per dataset in the original Tülu 2 dataset and our unfiltered version. Note that we deduplicate samples in the unfiltered mixture.
@@
WildChat (GPT-4 subset) (Zhao et al., 2024b) 100,000 235,028
@@
We provide a breakdown of the splits of the Tülu 2 SFT mixture and our ‘unfiltered’ version in Table 5. We provide a similar comparison between the Tülu 3 SFT mixture and our ‘unfiltered’ version in Table 6.
@@ G.2 RDS+ Weighted Pooling Details
Where $h_{i}$ is the last layer hidden state of the $i^{\text{th}}$ token.
Appendix H Visualization of Selected Samples
We visualize what samples get selected when selecting 326,000 samples using RDS+ from the Tülu 2 unfiltered pool in Figure 6, and what samples get selected when selecting with IFD, Top-PPL, random (unbalanced), and RDS+ from the Tülu 2 unfiltered pool in Figure 7.

### v2

@@ Abstract
Selecting high-quality training data from a larger pool is a crucial step when instruction-tuning language models, as carefully curated datasets often produce models that outperform those trained on much larger, noisier datasets. Automated data selection approaches for instruction-tuning are typically tested by selecting small datasets ($\sim$10k samples) from small pools ($\sim$200k samples). However, popular deployed instruction-tuned models often train on hundreds of thousands to millions of samples, subsampled from even larger data pools. We present a systematic study of how well data selection methods scale to these settings, selecting up to 2.5M samples from pools of up to 5.8M samples and evaluating across 7 diverse tasks. We show that many recently proposed methods fall short of random selection in this setting (while using more compute), and even decline in performance when given access to larger pools of data to select over. We find that a variant of representation-based data selection (RDS+), which uses weighted mean pooling of pretrained LM hidden states, consistently outperforms more complex methods across all settings tested – all whilst being more compute-efficient. Our findings highlight that the scaling properties of proposed automated selection methods should be more closely examined.
@@
Hamish Ivison1,2, Muru Zhang1,3, Faeze Brahman2, Pang Wei Koh1,2, Pradeep Dasigi2 1University of Washington, 2Allen Institute for AI, 3USC hamishiv@cs.washington.edu
@@ 1 Introduction
We test nine different data selection approaches in our setting, covering a variety of different approaches to data selection: gradient-based influence methods (Xia et al., 2024), embedding-based methods (Zhang et al., 2018; Hanawa et al., 2021; Ivison et al., 2023a), and loss/perplexity-based methods (Yin and Rush, 2024; Antonello et al., 2021; Marion et al., 2023; Ankner et al., 2024; Li et al., 2024b). Surprisingly (and contrary to prior work), we find that a simple embedding method we call RDS+ — using the hidden states of a pretrained LM — works best overall across all settings tested, consistently beating all other approaches. In particular, we find that:
@@
    Similarly, when selecting for multiple tasks at once, only 3 out of 7 selection methods beat random selection, with RDS+ performing best. We find that RDS+ even outperforms Tülu 2, which was trained on a human-curated mixture (§4.2).
@@
    Controlling for the compute used during selection, RDS+ outperforms random selection by an average of two points when selecting hundreds of thousands to millions of samples, whilst using less compute. This stands in contrast to prior work (Xia et al., 2024), which does not scale up to large enough sizes to observe the point at which RDS becomes more compute-efficient. This highlights that data selection techniques may only become particularly useful for larger-scale settings than those typically used in prior work. (§4.3)
We believe our findings highlight the importance of examining data selection methods at larger scales than considered in prior work, and show that some such methods can still provide benefits at these scales.
@@ 2 Related Work
Earlier approaches to data selection for instruction tuning relied primarily on careful human curation (Zhou et al., 2023) or extensive ablation studies to determine optimal data mixtures (Wang et al., 2023). Recently, many automated data selection approaches have emerged, ranging from using GPT-based quality assessment  (Chen et al., 2024; Wettig et al., 2024; Lu et al., 2023; Liu et al., 2024), embedding or perplexity features from model inference (Marion et al., 2023; Ivison et al., 2023a; Li et al., 2024b; Zhang et al., 2024c; Sachdeva et al., 2024; Li et al., 2024a; Du et al., 2023; Zhang et al., 2024b), gradient-based methods (Xia et al., 2024; Han et al., 2020; Zhang et al., 2024a; Yu et al., 2024), or shallower heuristics (Zhao et al., 2024a; Cao et al., 2024; Li et al., 2024d). These works typically focus on finding small subsets of larger datasets that can match or outperform using the entire dataset, often selecting subsets of roughly 10k samples in size (Chen et al., 2024; Xia et al., 2024; Li et al., 2024b) from datasets with 100s of thousands of samples. This does not match the known sizes of various instruction-tuning datasets used in practice, which range from 300k to millions of samples in size (Touvron et al., 2023; Llama Team, 2024; Ivison et al., 2023b; Lambert et al., 2024). In this work, we aim specifically to study how these methods perform when selecting large datasets from large pools, and find that many proposed methods fall short of random selection in this setting.
@@ 3 Data Selection Methods
Length.
We sort examples by length (in tokens) and take the longest samples. This has been shown to be a strong baseline by Zhao et al. (2024b), and is computationally cheap, only requiring computing the length of each sample in our data pool.
@@ 3.2 Data Pool
Figure 2: Size and makeup of data pools considered in this work (unfiltered Tulu 2, 3) and in past work (Xia et al., 2024; Chen et al., 2024; Li et al., 2024b). We provide the size of each pool on top of each bar. Each color represents a different dataset. See App. B for more details on data pool composition and exact per-source counts.
@@ 4.1 Single-Task Data Selection
Method MMLU GSM8k BBH TydiQA Codex Squad AlpacaEval Average Selecting 10k datapoints for single tasks from 200k data points Random (unbal.) 45.8 17.2 41.5 52.7 27.0 83.5 18.4 40.9 Random (bal.) 45.4 16.6 40.7 50.7 27.9 80.6 33.2 42.2 LESS 46.7 21.5 41.8 57.5 19.6 84.7 37.9 44.2 Embed (NV) 46.9 19.9 42.4 52.3 25.0 88.3 24.0 42.7 Embed (GTR) 47.1 20.2 42.4 52.4 27.7 88.3 16.7 42.1 Top-PPL 46.4 9.9 33.8 46.9 20.3 77.6 4.3 34.2 Mid-PPL 46.9 15.5 39.7 53.6 25.7 81.6 12.9 39.4 IFD 42.6 17.5 39.1 38.7 29.1 57.9 45.9 38.7 Length 43.2 14.2 40.2 54.3 30.4 82.3 40.0 43.5 RDS+ 47.0 20.9 42.7 52.3 31.8 88.2 42.3 46.4 Selecting 10k datapoints for single tasks from all 5.8M data points Random (unbal.) 46.9 17.4 42.1 53.2 26.8 83.6 15.1 40.7 Random (bal.) 45.3 16.5 40.2 50.9 28.2 78.0 29.0 41.1 Top-PPL 44.6 6.1 23.9 40.4 20.9 74.4 2.5 30.4 Mid-PPL 46.1 14.9 40.7 52.7 23.0 82.3 12.5 38.9 Embed (GTR) 45.0 29.9 42.8 45.5 29.7 88.3 43.8 46.4 Embed (NV) 47.0 23.1 41.1 51.2 29.1 88.8 10.9 41.6 IFD 41.9 13.0 37.9 40.0 26.4 46.6 44.6 35.8 Length 40.7 2.4 39.2 17.8 20.9 0.2 3.1 17.8 RDS+ 47.5 33.9 42.9 54.9 32.4 88.5 53.5 50.5
@@ 4.2 Multi-task Selection
Method MMLU GSM8k BBH TydiQA Codex Squad AlpacaEval Average Rand. (unbal) 52.2 18.0 44.5 55.3 25.7 81.5 33.9 44.5 Rand. (bal) 51.5 21.8 45.1 50.7 32.2 87.9 43.2 47.5 Top-PPL 49.1 10.5 39.4 49.4 21.6 80.3 5.6 36.6 Mid-PPL 53.1 13.3 42.8 52.4 20.3 86.2 20.7 41.3 Embed (GTR) 49.9 32.8 44.6 54.4 30.4 88.4 35.7 48.0 Embed (NV) 50.6 28.7 44.4 56.0 30.4 89.1 17.9 45.3 IFD 45.7 21.8 41.2 39.5 27.7 17.0 57.4 35.7 Length 50.0 16.4 38.8 54.9 25.0 89.3 58.5 47.6 Tülu 2 50.0 22.7 45.1 54.0 33.1 86.9 54.4 49.5 RDS+ 50.2 35.2 44.7 56.3 35.1 89.0 45.6 50.9 RDS+ - Wildchat 50.9 24.8 43.6 57.3 31.1 87.3 39.3 47.8 RDS+ - Arena Hard 48.1 36.2 43.9 51.8 31.8 81.3 59.4 50.4
@@
We explore using WildChat (Zhao et al., 2024c) and Arena Hard (Li et al., 2024c) as out of distribution query sets for selecting data. Crucially, this means we do not assume access to any query samples from our test tasks. We find that using Arena Hard samples performs close to RDS+, showing we can select high-quality samples without assuming any data from the evaluations in our test suite. This suggests that using a high-quality selection set results in a data mixture that generalizes well to unseen tasks.
@@ 4.3 Scaling Multi-task Selection
(Fig. 4). This shows that targeted selection is even useful when selecting datasets up to millions of data points in size. Furthermore, RDS+ performs similarly to training on all data (50.8 vs 50.9) even when selecting only 326k samples (roughly 6% of the overall pool), and outperforms training on all data when selecting over 1M samples.
@@
(Fig. 3). When taking into account the compute used for selection and training, we find that the extra compute incurred by RDS does not pay off until we select relatively large datasets (roughly $\geq 326$k), after which it becomes significantly better than random selection. This highlights that examining data selection performance at scale is important to show the potential of some methods.
@@ 5 Analysis
We examine qualitatively what samples get selected by RDS+, IFD, and PPL (see App. I for a visualization). For RDS+, We find that the sources selected vary depending on downstream evaluation, often with intuitive explanations: e.g., GSM8k requires strong chain-of-thought reasoning abilities, and so the chain-of-thought data is upweighted relative to the normal distribution. Similarly, for HumanEval-Codex the Code Alpaca dataset is upweighted.
@@ 6 Conclusion
In this work, we have explored how well a variety of data selection techniques work when selecting datasets of up to 2.5M samples from pools comprising of close to 6M data points. By examining data selection techniques in these settings, we find that many methods proposed in the past not only underperform random selection baselines, but also perform worse with larger pools of data. Embedding-based methods, and in particular RDS+, are the primary exception, outperforming all other data selection methods across various selection sizes. Finally, we find that using smaller models to select data can also work well, suggesting future work could further reduce the cost of RDS+ by exploring how best to use smaller proxy models. We believe that our results overall highlight the importance of testing data selection methods over large, diverse pools of data, and testing how well data selection approaches scale with respect to both data and compute. We will publicly release our data and code to aid in future work.
@@ Appendix B Data Pool Sources Breakdown
Source Tülu 2 Tülu 2 Unfil. LESS IFD Alpagasus FLAN V2 (Chung et al., 2022) 49,123 961,322 100,000 - - FLAN CoT (Chung et al., 2022) 49,747 398,439 100,000 - - Open Assist. (Köpf et al., 2023) 7,331 7,707 55,668 - - Dolly (Databricks, 2023) 0 15,007 15,011 - - GPT-4 Alpaca (Peng et al., 2023) 19,906 52,002 - - - Code Alpaca (Chaudhary, 2023) 20,016 20,022 - - - ShareGPT 111,912 100,054 - - - LIMA (Zhou et al., 2023) 1,018 1,030 - - - Wizard Evol-Instruct V2 (Xu et al., 2023) 29,810 142,802 - - - Open Orca (Lian et al., 2023) 29,683 4,111,858 - - - SciRIFF (Wadden et al., 2024) 7,468 7,535 - - - Alpaca (Taori et al., 2023) - - - 52,002 52,002 WizardLM (70K) (Xu et al., 2023) - - - 70,000 - Hardcoded 140 14 - - - Total 326,153 5,817,792 270,679 122,002 52,002
Table 5: Number of samples per dataset in the original Tülu 2 dataset and our unfiltered version alongside the data pools used in LESS (Xia et al., 2024), IFD (Li et al., 2024b), and Alpagasus (Chen et al., 2024). Note that we deduplicate samples in the Tülu 2 unfiltered mixture.
@@
WildChat (GPT-4 subset) (Zhao et al., 2024c) 100,000 235,028
@@
We provide a breakdown of the splits of the Tülu 2 SFT mixture and our ‘unfiltered’ version, along with other data pools from prior work, in Table 5. We provide a similar comparison between the Tülu 3 SFT mixture and our ‘unfiltered’ version in Table 6.
@@ G.2 RDS+ Weighted Pooling Details
Where $h_{i}$ is the last layer hidden state of the $i^{\text{th}}$ token. The rationale for this weighting is due to the causal mask of the decoder-only models we use for embedding: since the model is trained with a causal mask, the first input token does not see the rest of the sequence. Likewise, the second token does not see token 3 onwards, etc. As such, naive mean-pooling may bias the features towards earlier tokens in the sequence, since hidden states only accumulate features from previous tokens. The weighted mean-pooling strategy attempts to counter this by then weighting later tokens heavier.
Appendix H Visualization of RDS+
We provide a basic visualization of RDS+ in Figure 6.
Figure 6: Overview of RDS+ data selection. A pretrained LM encodes a data pool and a set of query points via weighted mean-pooling. We then compute cosine similarity between each point in the pool and the query set, and pick the top-K most similar points using Algorithm 1.
Appendix I Visualization of Selected Samples
We visualize what samples get selected when selecting 326,000 samples using RDS+ from the Tülu 2 unfiltered pool in Figure 7, and what samples get selected when selecting with IFD, Top-PPL, random (unbalanced), and RDS+ from the Tülu 2 unfiltered pool in Figure 8. We provide the raw per-source counts used to construct these figures in Tables 10 and 11 respectively.